OpenStudy (anonymous):

Complete the solution of the equation. Find the value of y when x equals -7. -3x+6y=-21

OpenStudy (anonymous):

how would you find the solution ?

OpenStudy (whpalmer4):

Find the value of \(y\) when \(x = -7\) How about you replace \(x\) with \((-7)\) in the equation? What does that give you?

OpenStudy (anonymous):

-3x(-7)+6y=-21?

OpenStudy (whpalmer4):

No, replace \(x\) with \((-7)\). You won't have any \(x\)'s present after you do that.

OpenStudy (anonymous):

Oh whoops my mistake so theres no x after the -7, -3(-7)+6y=-21 so you multiply ? -3*-7=21

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

mmm right so far I think? -3(-7)+6y=-21 is right... then -3*-7=21 is true so ^simplifies to 21 + 6y = -21 <-- solve for y

OpenStudy (anonymous):

subtract 21 from each side then divide by 6

OpenStudy (anonymous):

so 6y=-42 so y=-7

OpenStudy (anonymous):

subtract from which sides ?

OpenStudy (anonymous):

the side with the 6y on it

jigglypuff314 (jigglypuff314):

21 + 6y = -21 -21 -21 6y = -42

OpenStudy (anonymous):

then divide by 6 so y=-7

OpenStudy (anonymous):

ok help me out with this one bro Complete the solution of the equation. Findthe value of y when x equals 6. 4x+y=20

jigglypuff314 (jigglypuff314):

same idea :) plug in 6 for x to get 4(6) + y = 20 then solve for y

OpenStudy (anonymous):

24+y=20

OpenStudy (anonymous):

subtract 24 from the same side as y y=-4

OpenStudy (anonymous):

ok so first you 4(6)+y=20 24+y=20 ohhh ok i get your so the answer is y=-4

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

i thought you had to divide ?

jigglypuff314 (jigglypuff314):

only if/when there is a coefficient attached to the y like if it was "2y" you would divide by 2 but since it's just " y " you don't have to :)

OpenStudy (anonymous):

thanks guys this helped out big time !

jigglypuff314 (jigglypuff314):

glad we could help :)

OpenStudy (whpalmer4):

Sorry I had to turn into a pumpkin in the middle, glad @jigglypuff314 finished the problem for you!